OpenStudy (anonymous):

I am struggling with understanding the distributive properties of the problems I am doing. Can someone please help me step by step? y-2=-4/3(x+3)

OpenStudy (anonymous):

now multiply both sides by 3(x+3)

OpenStudy (anonymous):

what would u get?

OpenStudy (anonymous):

u there?

OpenStudy (anonymous):

The multiply fractions with whole numbers that messes me up every time! So -4/3(x+3) is hard for me.

OpenStudy (anonymous):

(y-2) ( x+3)3 = -4 3[y(x+3) -2(x+3)]=4

OpenStudy (anonymous):

That will get me the slop-intercept?

OpenStudy (anonymous):

wait..forget what i said

OpenStudy (anonymous):

The problem starts out (1,-3) m=-2/5

OpenStudy (anonymous):

I am suppose to graph the line passing through the given point and having the given slope. Give the slope-intercept form of the equation of the line if possible.

OpenStudy (anonymous):

so do (y+3) = -2(x-1)/5

OpenStudy (anonymous):

As far as I got was, y-3+3=-2/5x+2/5+3. Would that be right?

OpenStudy (anonymous):

Are you there?